Mathematics 242 – Principles of Analysis Definitions Quiz 5 – October 15, 2004

Your Name:

1. (15) What is the precise definition of the statement “the sequence (x_{n}) converges to the real number L?”

2. (10) What is the negation of your statement from part 1? (Express the negation in full symbolic for. “The sequence (x_{n}) does not converge to the real number L” is not an acceptable answer!)